5 January 2012 - 11.00AM
Salvors working on the Rena are preparing for severe bad weather expected to arrive on Saturday afternoon.
- Salvors are preparing for severe bad weather this weekend, with a low pressure system expected to arrive on Saturday afternoon, with the worst weather expected on Sunday. This is forecast to result in a maximum sea state of 6-7m, which is among the worst weather to be forecast so far.
- The weather for today and through until tomorrow is for cloudy periods and showers to continue, with swells of about 1m at Astrolabe Reef today.
- Salvors will continue to make the most of the good weather expected for today and tomorrow to remove containers. Fourteen containers were removed from the no.7 hold of the vessel yesterday and loaded onto the ST60, bringing the total number on board the barge to 15. This leaves an estimated 898 containers still on board Rena.
- Salvors will today focus on removing containers from the no.6 hold, as this presents the easiest access for container removal.
- Salvors will be working on pumping gas from the no.5 hold so to make the space safe for accessing containers once no.6 has been completed.
- Work is continuing on erecting a temporary gangway between the forward and aft sections of ship, with a more permanent structure to be built.
- Pumping of oil remnants from the no.5 starboard fuel tank was carried out for 35 minutes yesterday.
- The ship remains in a fragile but stable condition, with electronic sensors on board showing no significant change.
- Weather conditions yesterday meant divers were not able to inspect the underside of the vessel.
- Divers hope to be able to inspect the recent damage to the vessel today, but this will be dependent on the conditions and will only occur when it is safe for them to enter the water.
- Containers on the bow of the vessel continue to remain lashed down and those that are safely accessible have been fitted with transponders in anticipation of the severe weather front coming through at the weekend.
- The BOPRC Harbour master is also continuing to deal with a number of vessels and small craft breaching inside the exclusion zone, and reminds all water users that the 3NM zone remains in place around the Rena for operational safety reasons.
- A Braemar Howells team will this morning continue picking up container debris between Papamoa Beach and Maketu – comprising mainly plastic and milk powder, and some meat products.
- The company reports that a number of milk powder packets appear to have been opened after arriving on shore, and is urging members of the public to please not open or touch container materials, as this makes them more difficult to recover and is a potential health hazard. Anyone finding containers or container debris is asked to please report it.
- Searches by sonar vessels yesterday revealed 6 potential container-shaped targets, and attempts will be made to lower a drop camera to positively identify these further. Sonar searches will continue today.
- Reports yesterday from fishermen of wood in the water to the south west of Motiti Island are being followed up and any material will be recovered.
- A team will be checking the beach between Kaituna Cut and Maketu today for further container debris.
- 357 containers have now been landed, with none brought ashore yesterday.
Oil spill response
- Responders remain on standby and are ready to ramp up in anticipation of possible further releases of oil from Rena following the forecast bad weather this weekend.
- Shoreline assessment teams will today be inspecting beaches at Omanu and Papamoa to assess whether they can be signed off as clean.
- A team will be out at Matakana Island today assessing the area for operations next week.
- This morning’s over flight detected a 10x20m patch of oil with a metallic sheen at the port side of Rena, and a silvery sheen of oil about 10-30m wide and 3km long stretching from the vessel towards the north.
- A 10-person team is working on Leisure Island today to remove oil from rocky areas brought ashore by the recent weather.
- Contractors will be returning from the break and resuming work at various locations from Monday next week.
- Planned night operations at Rabbit Island were cancelled last night due to poor weather.
- 4 more birds have been successfully transferred to Massey University for care, bringing the total number of birds at Massey to 25.
- 6 of the birds at Massey are being prepared for release, most likely next week.
- The Motiti field team will continue monitoring the condition of wildlife there for the rest of this week.
- The wildlife facility at Te Maunga continues to be demobilised as less wildlife is being collected.
For further information contact:
Maritime New Zealand Media Line
Phone 04 499 7318
Back to media releases